1. Using the table, calculate the quantity of cars that need to be sold to avoid a loss at the net income level.

Selling price per car $20,000

Variable operating cost per car $18,000

Labor cost per car $500

Depreciation cost per car $1000

Fixed operating costs $10 million

Interest costs on debt financing $6 million

A. 5000 cars

B. 8000 cars

C. 20,000 cars

2. Which of the following is most likely to be a secondary source of liquidity for the firm?

A. Bank lines of credit

B. Excess cash in bank

C. Cash from sale of the firm's factory used for production
Answer #1

Solution 1:

Before we calculate the break even point as required in the question, please note that

  1. Depreciation is a part of fixed operating costs of a business and hence the 'depreciation cost per car' given is also included in the 'fixrd operating costs' of $10m and hence will not be considered to avoid double counting of depreciation expense.
  2. 'Labour cost per car' is already a part of 'Variable operating cost per car' and hence will not be considered to avoid double counting of expenses

Now, the number of cars that are required to be sold to avoid losses at net income level can be calculated using the below formula:-

Required sales (units)= (Fixed operating costs + interest costs)/Contribution per unit

Or,

Required sales (Units)= (Fixed operating costs + Interest costs)/(Selling price per car - variable cost per car')

Putting values from the question, we get as follows:

Required sales (Units)= (10,000,000 + 6,000,000)/(20,000 - 18,000)

Required Sales (Units)= 16,000,000/2,000 = 8,000 units

Therefore, 8,000 cars are required to be sold to avoid losses at net level and the correct option is B.

Solution 2:

Whenever a business requires financing or liquidity, the last source it would ideally want is through sale of any of the operating assets of the firm. If the firm sells its operating assets for liquidity, it hampers the existing business, impacts the brand, hampers company's image in the market and should be avoided.

The only time when a firm wants to sell its operating assets to get liquidity is if it wants to discontinue a segment or business and sells those segment assets in lieu of that. But under, normal circumstances, the firm would always prefer to acquire liquidity through bank credit or excess cash in bank instead of selling of its assets.

Therefore, in the given question, the correct option is C and Sale of firm's factory used for operations is most likely to be a secondary source of liquidity for the firm.